Suppose y varies directly as x and that y=21 when x=3. Find the value of y when x=5.
Answer by jojo14344(1513)

By Ratio and Proportion;
1st condition; x=3 and y=21
2nd condition; x=5 and y=?
so,
3/21 = 5/y
y = 5 * 21 /3
y = 35
It is just right to be y=35 (as shown by ratio above)as it varies directly to x.
When x goes up, so does y goes up!